OpenStudy (doglover1011010101101010101101):

A ferris wheel has a radius of 80 feet. Two particular cars are located such that the central angle between them is 165 degrees. to the nearest 10th, what is the length of the intercepted arc between those 2 cars on the ferris wheel ? show all work

OpenStudy (giantrobot):

\[160\pi \times\frac{ 165 }{ 360 }\] you get 230 feet.

OpenStudy (doglover1011010101101010101101):

thank you

OpenStudy (anonymous):

230 feet